OpenStudy (coolshubhs):

solve this as quadratic form : (x+1)(2x+3)(2x+5)(x+3)=945

OpenStudy (asnaseer):

I would first try to write 945 as a product of its prime factors.

OpenStudy (asnaseer):

can you do that?

OpenStudy (coolshubhs):

it is to be solved like (x+a)(x+b)(x+c)(x+d) where the sum of any two from a,b,c,d would be equal to other two...

OpenStudy (asnaseer):

you need to find 4 numbers that, when multiplied together, will give you 945 - correct?

OpenStudy (asnaseer):

those 4 numbers are: 1. x+1 2. 2x+3 3. 2x+5 4. x+3

OpenStudy (asnaseer):

in problems like this, it is usually best to start by breaking down 945 into its prime factors

OpenStudy (coolshubhs):

for example : (x+2)(x+3)(x+4)(x+1)=2 [(x+2)(x+3)][(x+4)(x+1)] =2 (in this group i have 2 product of square braces in which i have grouped 2+3 , 4+1 = 5 the same have to be done somehow in this question!

OpenStudy (asnaseer):

sorry but I do not understand what you are trying to achieve. I can solve this and find a value for x such that (x+1)(2x+3)(2x+5)(x+3)=945 by breaking 945 down into its prime factors. if you are trying to do something different, then I do not quite understand what it is that you are trying to do. maybe someone else will be able to help you better.

OpenStudy (coolshubhs):

okay can you plz do it your way?

OpenStudy (asnaseer):

I can certainly guide you to the solution. so first step is to break 945 into its prime factors. what do you get for that?

OpenStudy (coolshubhs):

3*3*3*5*7

OpenStudy (asnaseer):

good - now we need to group this into a product a 4 numbers - can you see how to group them?

OpenStudy (asnaseer):

*product of

OpenStudy (coolshubhs):

uhm like 3*3*3*5*7=9*3*5*7

OpenStudy (asnaseer):

yup - so the smallest of these is 3 - agreed?

OpenStudy (coolshubhs):

yup!

OpenStudy (asnaseer):

and the smallest in the product: (x+1)(2x+3)(2x+5)(x+3) is (x+1) so I would set: x + 1 = 3 and solve for x

OpenStudy (coolshubhs):

x=2 ?

OpenStudy (asnaseer):

yup! - now try substituting that into the equation and see if you get the right values

OpenStudy (coolshubhs):

i have three answers given in my text book : 2, -6 and one more value.. like complex no.

OpenStudy (asnaseer):

hmmm - I have never really come across a problem like this before and the way you described the solution above is not something I am familiar with. is there a specific name for this type of problem?

OpenStudy (asnaseer):

I'd be interested in learning about it.

OpenStudy (coolshubhs):

yup .. first yo have to convert them into quadratic form.. then solve.

OpenStudy (asnaseer):

so are you saying that you have to form the product of two quadratics, and then find all possible values for each one of those quads and then solve them?

OpenStudy (asnaseer):

is there a simpler example that you have solved that you can show me here - maybe then I will understand better?

OpenStudy (coolshubhs):

here's a similar question : http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.126457.html

OpenStudy (asnaseer):

wait - that uses the same method I used O.O

OpenStudy (coolshubhs):

but in my text book it has converted into quadratic form , then solved.

OpenStudy (coolshubhs):

can i show you the solution of this one.. ?

OpenStudy (asnaseer):

yes please - that may help to explain the process

OpenStudy (coolshubhs):

SOLUTION: (x+1)(x+2)(x+3)(x+4) = 120 [ (x+1)(x+4) ] [(x+2)(x+3)] =120 (x^2+5x+4) (x^2+5x+6)=120 put y=x^2+5x

OpenStudy (asnaseer):

oh I see now - let me see if I can try to solve this question in a similar fashion

OpenStudy (coolshubhs):

you can the two two constants in the square braces add to give the same value thats the idea... like in first brace (4+1=5) , in second (3+2=5). So that's the arrangement.

OpenStudy (coolshubhs):

*you can see

OpenStudy (asnaseer):

I believe you need to group them as follows: (x+1)(2x+3)(2x+5)(x+3) = (x+1)(x+3) * (2x+3)(2x+5)

OpenStudy (asnaseer):

then you'll get:\[(x^2+4x+3)(4x^2+16x+15)\]

OpenStudy (asnaseer):

and you can set: \(y=x^2+4x\)

OpenStudy (coolshubhs):

i have also given a hint for this ques. : LHS : 4[ (x+1) (x+3)] [ (x+3/2) (x+5/2) ]

OpenStudy (asnaseer):

so we end up with:\[(y+3)(4y+15)=945\]

OpenStudy (asnaseer):

and you can expand that and solve as a quadratic

OpenStudy (asnaseer):

\[4y^2+27y+45=945\]\[4y^2+27y-900=0\]\[(y-12)(4y+75)=0\]\[y=12\text{ or }-18\frac{3}{4}\]

OpenStudy (asnaseer):

so then you just use this to find x from the relation: \(y=x^2+4x\) does that seem correct to you?

OpenStudy (coolshubhs):

yes :D i got this thank you ... :D

OpenStudy (asnaseer):

thank YOU for teaching me a new technique! :)

OpenStudy (coolshubhs):

:D thank you too...